INFRARED SPECTRA OF HELIUM CLUSTERS

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

By observing vibration-rotation spectra of an infrared chromophore, X, the evolution of the physical properties of cold (∼0.2 K) helium clusters, ^{4}He_{N}-X, can be studied as a function of cluster size, N. The experiments involve direct absorption of a tunable infrared laser beam in a pulsed supersonic jet expansion, and cover a size range from N = 0 and 1 (the small molecule limit) up to N ∼ 100, often with atom-by-atom resolution. These results complement a remarkable body of data now available for very large clusters (“helium nanodroplets”) with N ∼ 10^{3} - 10^{4}. For the probe molecule, we have used OCS, N_{2}O, CO_{2}, CO, and (with limited results) SiH_{4}. A number of approaches are used for cluster size assignment, which is not given directly by experiment: dependence of the spectra on conditions such as jet backing pressure; comparison with microwave observations; and isotopic substitution. For OCS, which is also a favorite probe for helium nanodroplets, we resolve and assign spectra for virtually every N-value from 1 to 70, with microwave confirmation up to N = 39 [1]. Similar results are obtained for He_{N}-N_{2}O and He_{N}-CO_{2}, but some size ranges are incomplete due to spectral overlap. The He_{N}-CO spectra are quite different, but clusters can also be detected approaching N ∼ 100, with size assignments up to about 20. Notable results for all probes are that spectral lines remain sharp (∼0.001 cm^{−1}) even for larger clusters, and that B-values show broad oscillations as a function of cluster size, experimentally marking the evolution of superfluid helium solvation shells around the probe molecule.

[1] A.R.W. McKellar, Y. Xu, and W. Jäger, Phys. Rev. Lett. 97, 183401 (2006).